# State-insensitive wavelengths for light shifts and photon scattering from Zeeman states

### v1

@@ VII.2 Alkaline-earth ions
For singly-ionized alkaline-earth ions, the hyperfine splitting contributions are generally not dominated by the magnetic dipole moment. For many ions, the optimized detuning is thus less magic than for the majority of alkali atoms. Two exceptions are ${}^{43}$${}^{+}$ and ${}^{91}$${}^{+}$, where the optimized detuning should provide close to magic behavior. For many ions, the electric quadrupole contribution is actually larger than the magnetic dipole contribution. This is best exemplified by the two radium isotopes, where for some choices of $F$ and input polarization, the optimization procedure instead finds the trivial $\Delta\rightarrow\infty$ solutions rather than any local minima.
@@ VIII Applications
VIII.1 Magic wavelength for ground state Zeeman levels: state-independent ac Stark shift
@@
$\Delta E(F,m,\omega)=-\sum\limits_{\mu,\nu}\mathrm{Re}\left[\alpha^{mm}_{\mu\nu}(F,\omega)\right]E_{\mu}^{(-)}(\mathbf{r})E_{\nu}^{(+)}(\mathbf{r}).$ (30)
Unraveling the polarizability in terms of the dipole matrix elements, the shift reduces to
@@
The conditions we found above directly relate to the traditional conception of the ac Stark shift as scalar, vector, and tensor shifts. Our condition for preserved dipole ellipticity also makes the vector Stark shift zero, while the condition for equal dipole amplitudes removes the $m^{2}$ dependence of the tensor shift.
While the values for the optimized detuning we find here are too close to resonance to efficiently trap atoms without significant atomic saturation and heating, the magic detuning can be useful in other contexts. For instance, the state independent ac Stark shift is translated to state-insensitive mechanical forces, potentially allowing for optomechanical applications where the atomic internal state can be traced out. The magic detuning could then be used to explore driven-dissipative models for optomechanics in cavity QED [57, 58, 59].
@@ VIII.2 State-independent scattering in cavity QED
State-independent scattering was harnessed in Ref. [42], where a tweezer array of atoms was placed in a cavity to observe collective photon emission. Generically, this is a complicated problem involving atoms scattering drive photons into the cavity, exchanging photons with the cavity, and changing their internal state. However, by operating at the magic detuning, one can eliminate the internal atomic state, reducing the Hamiltonian to an effective drive of the cavity and a dispersive shift to its frequency. In this experiment, scattering at the magic detuning enabled the observation of superradiant and subradiant scattering from an atomic array into the cavity, and the reduction of the Hamiltonian to that of two-level systems was confirmed by polarization measurements of the output light. A full derivation of the Hamiltonian in this limit is provided in Appendix F.
@@ IX Conclusions
In conclusion, we have shown the generic existence of a “magic detuning” for scattering on the $D_{2}$ line of alkali atoms and alkaline-earth ions. At this detuning, a driven atom will respond in the same manner regardless of the magnetic sublevel the atom is in, to very good approximation. This means that the complex level structure of a single hyperfine level driven off-resonantly via three excited manifolds reduces to that of an off-resonantly driven two-level atom. This eliminates the need of magnetic state preparation in quantum optics and atomic physics experiments. At the magic detuning, a prepared superposition of magnetic states is not altered by the trapping light, nor does it decohere due to correlations with the scattered light.
@@
The magic detuning strictly occurs only in the absence of dc magnetic fields. Generally, Zeeman shifts must be small in comparison to the frequency range in which the response is close to magic. The magic condition also requires scattering to be via three excited hyperfine manifolds. Therefore, the nuclear spin has to be $I\geq 1$. The requirement for three excited manifolds means that scattering on a $D_{1}$ line cannot be magic. The necessary conditions for scattering on a $D_{1}$ line can be met independently, but they do not coincide (see Appendix E for details). Moreover, we have neglected scattering to other hyperfine ground levels assuming that the environment is frequency selective (such as, for instance, a cavity). If the environment is not frequency selective, such as in free space, the lifetime of atoms would be limited by the rate of spontaneous emission to the other hyperfine level, as spontaneous emission to both hyperfine levels cannot be state-independent at the same detuning. Repopulation of the desired hyperfine level is possible through a repump field, which can be trivially magic if the hyperfine splitting between the two hyperfine levels is much larger than that between the excited levels, such that the pump is effectively infinitely detuned. However, spontaneous emission and repumping will still decohere the atomic state. This limits the usefulness of the magic detuning in free space to situations where strictly maintaining coherence between the magnetic states is not crucial.
@@
The level to which the magic detuning conditions hold depends on the origin of the hyperfine splitting of the excited states. If the magnetic dipole dominates over the electric quadrupole (and higher order moments), the approximation is very good. While we have focused on single-valence electron atoms and ions, the same physics may hold for the multi-electron case. To find Zeeman-insensitive frequencies, one must find a combination of the quantum numbers that meets the necessary criterion given in Eq. (56) in App. C. One has to be careful with multi-valence electron atoms, where the spectrum is more cluttered and hyperfine splitting is more complicated as it must include “configuration interactions” [45, 63]. However, as evidenced by the case of ${}^{85}$b, where the electric quadrupole contribution to the hyperfine splitting is of similar magnitude to that of the magnetic dipole, the magic detuning is very robust to other sources of hyperfine splitting. Therefore, if the hyperfine splitting is still given by the magnetic dipole contribution to even a broad approximation, the two-level picture may still be a good approximation. The presence of a magic detuning may also be interesting in the context of molecules, where magic conditions for trapping have been found for rotational states [64, 65]. However, the presence of a cluttered spectrum due to rotational and vibrational states may prove problematic.
@@ Appendix A Evaluating the dipole matrix elements
$r_{F,F^{\prime}}^{m}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&1&-m-1\end{pmatrix},$ (33e)
$s_{F,F^{\prime}}^{m}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&-1&-m+1\end{pmatrix},$ (33j)
$t_{F,F^{\prime}}^{m}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&0&-m\end{pmatrix},$ (33o)
@@ A.1 Expressions for Wigner 3-j symbols
$\begin{pmatrix}F&1&F^{\prime}\\ m&1&-m-1\end{pmatrix}=(-1)^{2F+m-F^{\prime}+1}\sqrt{\frac{2(F+F^{\prime}-1)!}{(F^{\prime}-F+1)!(F-F^{\prime}+1)!(F+F^{\prime}+2)!}\frac{(F-m)!(F^{\prime}+m+1)!}{(F+m)!(F^{\prime}-m-1)!}},$ (34c)
$\begin{pmatrix}F&1&F^{\prime}\\ m&-1&-m+1\end{pmatrix}=(-1)^{F+m}\sqrt{\frac{2(F+F^{\prime}-1)!}{(F^{\prime}-F+1)!(F-F^{\prime}+1)!(F+F^{\prime}+2)!}\frac{(F+m)!(F^{\prime}-m+1)!}{(F-m)!(F^{\prime}+m-1)!}}.$ (34f)
@@
$\begin{pmatrix}F&1&F+1\\ m&0&-m\end{pmatrix}=(-1)^{F-1+m}\sqrt{\frac{2(F+1-m)(F+1+m)}{(2F+1)(2F+2)(2F+3)}},$ (35c)
$\begin{pmatrix}F&1&F+1\\ m&0&-m\end{pmatrix}=(-1)^{F-1+m}\frac{2m}{\sqrt{2F(2F+1)(2F+2)},}$ (35f)
$\begin{pmatrix}F&1&F-1\\ m&0&-m\end{pmatrix}=(-1)^{F+m}\sqrt{\frac{2(F+m)(F-m)}{(2F-1)2F(2F+1)}}.$ (35i)
@@ A.2 Expressions for Wigner 6-j symbols
$\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&1\\ F^{\prime}&F&I\end{Bmatrix}$ $=\left(-1\right)^{F^{\prime}+F+1}\sqrt{\frac{\left(F^{\prime}-F+1\right)!\left(F-F^{\prime}+1\right)!\left(F^{\prime}+F+2\right)\left(F^{\prime}+F+1\right)\left(F^{\prime}+F\right)}{6\left(F+I+\frac{3}{2}\right)\left(F+I+\frac{1}{2}\right)\left(F^{\prime}+I+\frac{3}{2}\right)\left(F^{\prime}+I+\frac{1}{2}\right)}}$ (36c)
$\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F^{\prime}&F&I\end{Bmatrix}$ $=\left(-1\right)^{F^{\prime}+I+3/2}\left(I+F^{\prime}+\frac{5}{2}\right)!$ (36f)
×$\times\sqrt{\frac{\left(F^{\prime}+\frac{3}{2}-I\right)!\left(I-F+\frac{3}{2}\right)!\left(F^{\prime}+F-1\right)!}{12\left(I+F+\frac{3}{2}\right)\left(I+F+\frac{1}{2}\right)\left(F^{\prime}+I-\frac{3}{2}\right)!\left(F^{\prime}-F+1\right)!\left(F-F^{\prime}+1\right)!\left(F+I+\frac{5}{2}\right)!\left(F^{\prime}+F+2\right)!}}$ (36g)
@@ Appendix B Derivation of $\Delta_{\perp}$
$a$ $=2F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}+(1-2F)(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (42g)
$b$ $=\zeta_{-1}2F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(\zeta_{-1}+\zeta_{+1})(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}+\zeta_{+1}(1-2F)(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (42n)
$c$ $=-(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}\zeta_{-1}\zeta_{+1}.$ (42q)
@@
$A$ $=2F(2F+3)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2},$ (43c)
$B$ $=-(4F+2)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2},$ (43f)
$C$ $=(1-2F)(2F+2)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (43i)
such that the coefficients of the quadratic equation read
@@
This cannot be true for any combination of $F$ and $I$ that satisfy $F=I\pm J$ and $F\geq 1$, and so there are always two roots to the quadratic equation. For those roots to be real we require
@@ Appendix C Derivation of $\Delta_{\parallel}$
C.1 Proof that there is only one root to the equation for $\Delta_{\parallel}$ for $J^{\prime}=J+1$
@@ D.1 $\Delta_{\perp}$ when the hyperfine splitting is dominated by the magnetic dipole contribution
Using the combination of Wigner 6-j coefficients in Eq. (56) to eliminate the $F^{\prime}=F-1$ terms, we arrive at the following form for the quadratic coefficients in Eq. (B):
$\frac{a}{A_{\mathrm{hfs}}}$ $=4F(2F+1)\left[\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}-\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}\right],$ (66e)
$\frac{b}{A_{\mathrm{hfs}}}$ $=2F(2F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}+2F(2F+1)^{2}\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2},$ (66j)
$\frac{c}{A_{\mathrm{hfs}}}$ $=F(4F+2)(F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}.$ (66m)
@@
$\frac{\Delta_{\perp}}{A_{\mathrm{hfs}}}$ $=\frac{-(F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}}{\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}-\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}},$ (69g)
$\frac{\Delta_{\perp}}{A_{\mathrm{hfs}}}$ $=\frac{1}{2}.$ (69h)

### v2

@@ VII.2 Alkaline-earth ions
For singly-ionized alkaline-earth ions, the hyperfine splitting contributions are generally not dominated by the magnetic dipole moment. For many ions, the optimized detuning is thus less magic than for the majority of alkali atoms, as shown in Table 2. Two exceptions are 43Ca+ and 91Sr+, where the optimized detuning should provide close to magic behavior. For many ions, the electric quadrupole contribution is actually larger than the magnetic dipole contribution. This is best exemplified by the two radium isotopes, where for some choices of $F$ and input polarization, the optimization procedure instead finds the trivial $\Delta\rightarrow\infty$ solutions rather than any local minima.
@@ VIII Applications
VIII.1 Magic wavelength for hyperfine manifolds: state-independent ac Stark shift
@@
$\Delta E(F,m,\omega)=-\sum\limits_{\mu,\nu}\mathrm{Re}\left[\alpha^{mm}_{\mu\nu}(F,\omega)\right]E_{\mu}^{(-)}(\mathbf{r})E_{\nu}^{(+)}(\mathbf{r}),$ (30)
where $E_{\mu}^{(-)}(\mathbf{r})=\braket{\hat{E}_{\mu}^{(-)}(\mathbf{r})}$ is the expectation value of the field. Unraveling the polarizability in terms of the dipole matrix elements, the shift reduces to
@@
The conditions we found above directly relate to the traditional description of the ac Stark shift in terms of scalar, vector, and tensor shifts. Our condition for preserved dipole ellipticity makes the vector Stark shift zero, while the condition for equal dipole amplitudes removes the $m^{2}$ dependence of the tensor shift.
While the values for the optimized detuning we find here are too close to resonance to efficiently trap atoms without significant atomic saturation and heating, the magic detuning can be useful in other contexts. For instance, the state independent ac Stark shift is translated to state-insensitive mechanical forces, potentially allowing for optomechanical applications where the Zeeman state can be traced out. The magic detuning could then be used to explore optomechanical driven-dissipative phase transitions [68, 69, 70].
@@ VIII.2 State-independent scattering in cavity QED
Our predictions for state-independent light scattering were experimentally confirmed in Ref. [52]. In this experiment, a tweezer array of up to eight single rubidium atoms were placed in a cavity and driven with linearly polarized light. The atoms were able to scatter light into two degenerate cavity modes with polarization parallel or perpendicular to the initial drive light. Generically, this is a complicated problem involving atoms scattering drive photons into both polarization modes of the cavity, exchanging photons with the cavity, and changing their Zeeman state in the process. However, by operating at the magic detuning, one can eliminate the Zeeman state, reducing the Hamiltonian to an effective drive of the cavity and a dispersive shift to its frequency. In this experiment, scattering at the magic detuning enabled the observation of superradiant and subradiant scattering from an atomic array into the cavity. Operation at the magic detuning guarantees that the induced atomic dipoles have the same polarization as the drive light and that only Rayleigh scattering occurs. The former was confirmed by measurements of the polarization of the cavity output light, while the latter was confirmed by the scaling of the scattered field with atom number. A full derivation of the Hamiltonian in this limit is provided in Appendix F.
@@ IX Conclusions
In conclusion, we have shown the generic existence of a “magic detuning” for scattering on the $D_{2}$ line of alkali atoms and alkaline-earth ions. At this detuning, a driven atom will respond in the same manner regardless of its Zeeman state, to very good approximation. Therefore, the complex level structure of a single hyperfine manifold driven off-resonantly via three excited manifolds reduces to that of an off-resonantly driven two-level atom. This eliminates the need of Zeeman state preparation in quantum optics and atomic physics experiments. At the magic detuning, a prepared superposition of Zeeman states is not altered by the trapping light, nor does it decohere due to correlations with the scattered light.
@@
The magic detuning strictly occurs only in the absence of dc magnetic fields. Generally, Zeeman shifts must be small in comparison to the frequency range in which the response is close to magic. The magic condition also requires scattering to be via three excited hyperfine manifolds. Therefore, the nuclear spin has to be $I\geq 1$. Scattering via only two excited hyperfine manifolds cannot be magic. The necessary conditions for scattering on such a $D_{1}$ type line can be met independently, but they do not coincide (see Appendix E for details). Moreover, we have neglected scattering to other hyperfine ground levels assuming that the environment is frequency selective (such as, for instance, a cavity). If the environment is not frequency selective, such as in free space, the lifetime of atoms would be limited by the rate of spontaneous emission to the other ground state hyperfine manifold, as spontaneous emission to both hyperfine manifolds cannot be state-independent at the same detuning. Repopulation of the desired hyperfine level is possible through a repump field, which can be trivially magic if the hyperfine splitting between the two hyperfine levels is much larger than that between the excited levels, such that the pump is effectively infinitely detuned. However, spontaneous emission and repumping will still decohere the atomic state. This limits the usefulness of the magic detuning in free space to situations where strictly maintaining coherence between the Zeeman states is not crucial.
@@
The level to which the magic detuning conditions hold depends on the origin of the hyperfine splitting of the excited states. If the magnetic dipole dominates over the electric quadrupole (and higher order moments), the approximation is very good. While we have focused on single-valence electron atoms and ions, the same physics may hold for the multi-electron case. To find state-insensitive frequencies for Zeeman states, one must find a combination of the quantum numbers that meets the necessary criterion given in Eq. (54) in App. C. One has to be careful with multi-valence electron atoms, where the spectrum is more cluttered and hyperfine splitting is more complicated as it must include configuration interactions [56, 74]. However, as evidenced by the case of 85Rb, where the electric quadrupole contribution to the hyperfine splitting is of similar magnitude to that of the magnetic dipole, the magic detuning is very robust to other sources of hyperfine splitting. Therefore, if the hyperfine splitting is still given by the magnetic dipole contribution to even a broad approximation, the two-level picture may still hold. The presence of a magic detuning may also be interesting in the context of molecules [75], where magic conditions for trapping have been found for rotational states [76, 77]. However, the presence of a cluttered spectrum due to rotational and vibrational states may prove problematic.
@@ Appendix A Evaluating the dipole matrix elements
$r_{m,F,F^{\prime}}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&1&-m-1\end{pmatrix},$ (33a)
$s_{m,F,F^{\prime}}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&-1&-m+1\end{pmatrix},$ (33b)
$t_{m,F,F^{\prime}}$ $=(-1)^{m}\sqrt{2F^{\prime}+1}\begin{Bmatrix}J&J^{\prime}&1\\ F^{\prime}&F&I\end{Bmatrix}\begin{pmatrix}F&1&F^{\prime}\\ m&0&-m\end{pmatrix},$ (33c)
@@ A.1 Expressions for Wigner 3-j symbols
$\begin{pmatrix}F&1&F^{\prime}\\ m&1&-m-1\end{pmatrix}=(-1)^{2F+m-F^{\prime}+1}\sqrt{\frac{2(F+F^{\prime}-1)!}{(F^{\prime}-F+1)!(F-F^{\prime}+1)!(F+F^{\prime}+2)!}\frac{(F-m)!(F^{\prime}+m+1)!}{(F+m)!(F^{\prime}-m-1)!}},$ (34a)
$\begin{pmatrix}F&1&F^{\prime}\\ m&-1&-m+1\end{pmatrix}=(-1)^{F+m}\sqrt{\frac{2(F+F^{\prime}-1)!}{(F^{\prime}-F+1)!(F-F^{\prime}+1)!(F+F^{\prime}+2)!}\frac{(F+m)!(F^{\prime}-m+1)!}{(F-m)!(F^{\prime}+m-1)!}}.$ (34b)
@@
$\begin{pmatrix}F&1&F+1\\ m&0&-m\end{pmatrix}=(-1)^{F-1+m}\sqrt{\frac{2(F+1-m)(F+1+m)}{(2F+1)(2F+2)(2F+3)}},$ (35a)
$\begin{pmatrix}F&1&F+1\\ m&0&-m\end{pmatrix}=(-1)^{F-1+m}\frac{2m}{\sqrt{2F(2F+1)(2F+2)},}$ (35b)
$\begin{pmatrix}F&1&F-1\\ m&0&-m\end{pmatrix}=(-1)^{F+m}\sqrt{\frac{2(F+m)(F-m)}{(2F-1)2F(2F+1)}}.$ (35c)
@@ A.2 Expressions for Wigner 6-j symbols
$\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&1\\ F^{\prime}&F&I\end{Bmatrix}$ $=\left(-1\right)^{F^{\prime}+F+1}\sqrt{\frac{\left(F^{\prime}-F+1\right)!\left(F-F^{\prime}+1\right)!\left(F^{\prime}+F+2\right)\left(F^{\prime}+F+1\right)\left(F^{\prime}+F\right)}{6\left(F+I+\frac{3}{2}\right)\left(F+I+\frac{1}{2}\right)\left(F^{\prime}+I+\frac{3}{2}\right)\left(F^{\prime}+I+\frac{1}{2}\right)}}$ (36a)
$\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F^{\prime}&F&I\end{Bmatrix}$ $=\left(-1\right)^{F^{\prime}+I+3/2}\left(I+F^{\prime}+\frac{5}{2}\right)!$
×$\times\sqrt{\frac{\left(F^{\prime}+\frac{3}{2}-I\right)!\left(I-F+\frac{3}{2}\right)!\left(F^{\prime}+F-1\right)!}{12\left(I+F+\frac{3}{2}\right)\left(I+F+\frac{1}{2}\right)\left(F^{\prime}+I-\frac{3}{2}\right)!\left(F^{\prime}-F+1\right)!\left(F-F^{\prime}+1\right)!\left(F+I+\frac{5}{2}\right)!\left(F^{\prime}+F+2\right)!}}$ (36b)
@@ Appendix B Derivation of $\Delta_{\perp}$
$a$ $=2F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}+(1-2F)(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (42a)
$b$ $=\zeta_{-1}2F(2F+3)\begin{Bmatrix}J&J^{\prime}&1\\ F+1&F&I\end{Bmatrix}^{2}-(\zeta_{-1}+\zeta_{+1})(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}+\zeta_{+1}(1-2F)(2F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (42b)
$c$ $=-(4F+2)\begin{Bmatrix}J&J^{\prime}&1\\ F&F&I\end{Bmatrix}^{2}\zeta_{-1}\zeta_{+1}.$ (42c)
@@
$A$ $=2F(2F+3)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2},$ (43a)
$B$ $=-(4F+2)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2},$ (43b)
$C$ $=(1-2F)(2F+2)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F-1&F&I\end{Bmatrix}^{2},$ (43c)
such that the coefficients of the quadratic equation read as
@@
This cannot be true for any combination of $F$ and $I$ that satisfies $F=I\pm J$ and $F\geq 1$, and so there are always two roots to the quadratic equation. For those roots to be real we require
@@ Appendix C Derivation of $\Delta_{\parallel}$
C.1 Proof that there is only one root to the equation for $\Delta_{\parallel}$ for scattering on the $D_{2}$ line
@@ D.1 $\Delta_{\perp}$ when the hyperfine splitting is dominated by the magnetic dipole contribution
Using the combination of Wigner 6-j coefficients in Eq. (54) to eliminate the $F^{\prime}=F-1$ terms, we arrive at the following form for the quadratic coefficients in Eq. (42):
$\frac{a}{A_{\mathrm{hfs}}}$ $=4F(2F+1)\left[\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}-\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}\right],$ (64a)
$\frac{b}{A_{\mathrm{hfs}}}$ $=2F(2F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}+2F(2F+1)^{2}\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2},$ (64b)
$\frac{c}{A_{\mathrm{hfs}}}$ $=F(4F+2)(F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}.$ (64c)
@@
$\frac{\Delta_{\perp}}{A_{\mathrm{hfs}}}$ $=\frac{-(F+1)\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}}{\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F+1&F&I\end{Bmatrix}^{2}-\begin{Bmatrix}\frac{1}{2}&\frac{3}{2}&1\\ F&F&I\end{Bmatrix}^{2}},$ (67a)
$\frac{\Delta_{\perp}}{A_{\mathrm{hfs}}}$ $=\frac{1}{2}.$ (67b)